OpenStudy (anonymous):

Simplify (4x2 - 2x + 5) + (2x2 - 2x - 8) a: 2x2 + 13 b: 6x2 - 4x - 3 c: 6x2 - 4x + 13 d: 6x2 - 3

OpenStudy (janu16):

B

OpenStudy (anonymous):

thanks

OpenStudy (janu16):

Np

OpenStudy (mrnood):

look at the first and last terms in each bracket add the first pair that will give you the first term in the answer add the last pair - that will give you the last term in the answer there is only one answer that works @Janu16 the poster cannot know if you are right if you just post the answer @sistrunk23 I suggest you check the answer - it is quite likely wrong

OpenStudy (janu16):

Step 1 : Raise x to the 2nd power Exponentiation : Equation at the end of step 1 : (((4•(x2))-2x)+5)+(((2•x2)-2x)-8) Step 2 : Raise x to the 2nd power Exponentiation : Equation at the end of step 2 : (((4 • x2) - 2x) + 5) + (2x2 - 2x - 8) Step 3 : Simplify 4x2-2x+5 + 2x2-2x-8 Trying to factor by splitting the middle term 3.1 Factoring 6x2-4x-3 The first term is, 6x2 its coefficient is 6 . The middle term is, -4x its coefficient is -4 . The last term, "the constant", is -3 Step-1 : Multiply the coefficient of the first term by the constant 6 • -3 = -18 Step-2 : Find two factors of -18 whose sum equals the coefficient of the middle term, which is -4 . -18 + 1 = -17 -9 + 2 = -7 -6 + 3 = -3 -3 + 6 = 3 -2 + 9 = 7 -1 + 18 = 17 Observation : No two such factors can be found !! Conclusion : Trinomial can not be factored Final result : 6x2 - 4x - 3

OpenStudy (janu16):

Im correct